Circular No. 8895

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://www.cfa.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

NOVA PUPPIS 2007 Alfredo Jose Serra Pereira, Carnaxide, Portugal, reports his discovery

of an apparent nova (mag 7.0) during his regular visual patrol on Nov. 14.23 UT with 14×100 binoculars, the position given as α = 8^{h}16^{m}.2, δ = −34^{o}15′ (equinox 2000.0); he had nova-hunted for 625.85 hr since his discovery of the 2001 nova V4740 Sgr (cf. IAUC 7706). Pereira adds that nothing was detected at this position in his visual searches during Nov. 6.23, 7.22, 8.23, 10.23, and 11.22 down to mag ∼ 8; additional monitoring during Nov. 14.243–14.256 also yielded mag 7.0. Following requests by the Central Bureau, CCD astrometry for the variable has been reported by J. E. McGaha (Tucson, AZ, U.S.A., 0.62-m f/5.1 reflector) and by J. Young and H. Rhoades (Table Mountain 0.61-m f/16 Cassegrain reflector). From his 1-s exposure on Nov. 14.411, McGaha measures α = 8^{h}16^{m}18^{s}.01, δ = −34^{o}15′24′′.1, red mag 6.5 (USNO-A2.0 reference-star catalogue). From four exposures “in surprisingly good seeing for the low altitude” on Nov. 14.47–14.48, Young measures average position end figures 17^{s}.99, 25′′.0, mag 6.7 in a red filter (USNO-B1.0 reference-star catalogue); Young also estimated visual mag 6.4 at this time. Young adds that comparison with a red image from the Digitized Sky Survey shows only an extremely faint star (mag ∼ 20) at the same spot as the apparent nova (located in a fairly empty area surrounded by “a busy star field”).

COMETS C/2007 Q8, C/2007 Q9, AND C/2007 R6–R8 (SOHO) Further to IAUC 8893, additional Kreutz-sungrazing comets have been

found on SOHO website images, their “discovery” observations tabulated below. C/2007 Q8, also found by R. Kracht, was bright (mag ∼ 4.5) and elongated. C/2007 Q9, which was also clearly visible in SECCHI HI-1B images for several hours, was also quite bright (mag ∼ 5.5) and slightly elongated. C/2007 R6 was very faint (mag ∼ 8) and slightly diffuse. C/2007 R7 and C/2007 R8 were stellar in appearance and faint (mag ∼ 7) in C3 images; C/2007 R7 appeared small, diffuse, and very faint in C2 images.

Comet 2007 UT α_{2000} δ_{2000} Inst. F MPEC
C/2007 Q8 Aug. 28.596 10^{h}05.^{m}2 + 8^{o}30′ C3  HS 2007-U35
C/2007 Q9 28.988 10 07.2 + 8 26 C3 JS 2007-U35
C/2007 R6 Sept. 1.075 10 31.5 + 8 00 C2  HS 2007-U35
C/2007 R7  8.321 10 52.2 + 4 09 C3/2 HS 2007-U35
C/2007 R8  9.863 10 53.8 + 4 32 C3  BZ 2007-U36
2007 November 14  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2007 CBAT  Daniel W. E. Green